OpenStudy (lukecrayonz):

For f(x)=-x^2-4x-1, find f(x+h)-f(x)/h. I keep doing it but I'm getting the wrong answer, any help?:)

OpenStudy (anonymous):

go slow

OpenStudy (anonymous):

first find \(f(x+h)\)

OpenStudy (debbieg):

Is this what you're starting with for f(x+h)? \[f(x+h)=(x+h)^2-4(x+h)-1\]

OpenStudy (debbieg):

Oops, sorry, I missed the negative sign out in front...

OpenStudy (debbieg):

\[f(x+h)=-(x+h)^2-4(x+h)-1\]

OpenStudy (lukecrayonz):

@satellite73 for f(x+h) I get: -x^2-2hx-4x-h^2-4h-1

OpenStudy (anonymous):

looks good now subtract \(f(x)\)

OpenStudy (anonymous):

everything without an \(h\) should be gone when you are done